# Flow-Induced Phase Separation for Active Brownian Particles in Four-Roll-Mill Flow

### 28 Oct 2025 v1

@@ Abstract
We investigate the collective dynamics of active Brownian particles (ABPs) subjected to a steady two-dimensional four-roll-mill flow using numerical simulations. By varying the packing fraction ($\phi$), we uncover a novel flow-induced phase separation (FIPS) that emerges beyond a critical density ($\phi\geq 0.6$). The mean-square displacement (MSD) exhibits an intermediate bump between ballistic and diffusive regimes, indicating transient trapping and flow-guided clustering. The effective diffusivity decreases quadratically with $\phi$, while the drift velocity remains nearly constant, demonstrating that large-scale transport is primarily dictated by the background flow. Number fluctuations show a crossover from normal to giant scaling, signaling the onset of long-range density inhomogeneities in the FIPS regime. Our findings provide new insights into the coupling between activity, crowding, and flow, offering a unified framework for understanding phase behavior in driven active matter systems.
@@ I Introduction
The study of self-propelled particle systems gained significant attention following the pioneering work of Vicsek et al. [1], which introduced a minimal model exhibiting collective motion in the absence of aligning fields. This line of research revealed that the motion of active particles is not only governed by their internal propulsion mechanisms, but is also strongly influenced by the surrounding environment, including background flow fields, confinement, or hydrodynamic interactions [2, 3, 4, 5]. With the rapid advancement of computational resources in recent years, active matter has become a thriving area of research offering deep insights into nonequilibrium statistical mechanics.
@@
While the active Brownian particle (ABP) model neglects hydrodynamic, it has proven to be an indispensable minimal model that captures many features of active matter [14, 7, 19]. In ABPs, each particle self-propels at a constant speed and undergoes rotational diffusion, breaking detailed balance, and giving rise to non-equilibrium behavior even in simple settings [14, 20].
In real world scenarios, active particles are not isolated but are often immersed in complex fluid environments or interact with background flows, due to external fields, confinement [21, 22, 23]. The interaction between active particles and background flows can give rise to complex dynamical behaviors that are not captured only by considering the ABP model in the absence of external flow influences. In our recent work [20], we studied ABPs in a two-dimensional steady flow field and discovered a new non-equilibrium phase, which we term flow-induced phase separation (FIPS), driven by the ratio of persistence length of ABPs and background flow.
In this study, we investigate the onset and characteristics of the FIPS phase by systematically varying the packing fraction of ABPs in a two-dimensional background flow [20]. As MIPS is known to be strongly dependent on packing fraction [14, 17, 24], we explore FIPS by varying the packing fractions, which range from $0.1$ to $0.9$ to understand how packing fraction influences clustering and phase behavior under flow influence.
To characterize the system’s behavior across this range of packing fraction, we compute several key observables: number fluctuations[14, 25, 26, 27], mean squared displacement (MSD) [20, 14, 10], cluster size distributions (CSD) [28, 20], and Probability distribtuin function(PDF) of Okubo-Weiss parameter ($\Lambda$) [29, 30].
@@ II Model and Simulations
This flow field yields a periodic vorticity pattern in the square simulation domain of size $0\penalty 10000\ L]^{2}$, with $L=2\pi$, consisting of a pair of clockwise and counterclockwise vortices, separated by narrow regions dominated by strain [20, 31]. This flow has been investigated both experimentally and theoretically in the context of passive tracer mixing [32, 33]. We have set the amplitude of the velocity field to $\rm U_{0}=0.5$.
We use the Euler–Maruyama method to numerically integrate equations (1a) and (1b). The parameters are fixed as $\rm\mu_{t}=100$, $\rm k=1$, and $\rm v_{p}=0.2$. Based on the characteristic velocity scales of the ABPs and the background flow, we define a dimensionless scaled speed $\rm V=\frac{v_{p}}{U_{0}}$} [20]. A characteristic timescale for ABPs is given by $\rm\nu_{r}^{-1}$, while that for the background flow can be defined from the characteristic vorticity $\rm\omega_{0}=2U_{0}$. The dimensionless scaled time is then defined as the ratio of these two timescales, $\tau=\omega_{0}/(2\pi\nu_{r})$. In all simulations, we set $\rm V=0.4$} and $\tau\sim 50$. Based on our experience from prior work, we have chosen this particular $V$ and $\tau$ such that for $\phi=0.7$, we will get a flow-induced phase separation (FIPS), where ABPs try to phase separate in the strain dominated region of the flow [20]. We neglect the back-reaction of particles on the background flow, i.e., the flow field remains unaffected by the presence of ABPs [20]. The particle radius is fixed at $\rm a=0.02$. The packing fraction is defined as $\rm\phi=\frac{N\pi a^{2}}{L^{2}}$ and is varied from $0.1$ to $0.8$ by changing the number of particles $\rm N$ (e.g., $\rm N=21991$ for $\phi=0.7$). All simulations were performed for $10^{4}$ time units, which is sufficient to fully capture the diffusive regime. In all simulations, $\phi$ serves as the control parameter with $\phi\in[0.1,0.9]$.
@@ III Results
In the following sub-sections, we first make a qualitative comparison between MIPS and FIPS across varying $\phi$. We then present the mean-square displacement (MSD) to estimate drift velocity and diffusivity, followed by analyses of number fluctuations and cluster size distributions to quantify heterogeneity. Finally, to characterize the dependence of FIPS on flow topology, we examine the Okubo–Weiss parameter distribution.
@@ III.1 Flow-Induced Phase Separation
At high packing fraction ($\phi=0.7$), both systems exhibit phase separation, though their morphologies differ qualitatively. Without flow, we observe standard MIPS, where dense clusters form spontaneously and diffuse throughout the domain (Fig. 1c). In contrast, under imposed flow, the dense phase becomes localized in the strain-dominated regions, forming a characteristic four-lobed structure corresponding to the vortex cores of the four-roll mill flow (Fig. 1f). A systematic variation of $\phi$ from ($\rm\phi\in[0.1,0.9]$) is shown in Appendix Fig. 7, demonstrating that even at very high packing fractions ($\phi=0.9$), the four distinct low-density regions persist, coinciding with the four vortex cores of the background flow.
@@ III.2 Mean Square Displacement (MSD)
$\rm\langle[\Delta\boldsymbol{r}(t)]^{2}\rangle=4D_{e}\left[t+\frac{1}{\nu_{r}}(e^{-\nu_{r}t}-1)\right],$ (5)
@@
where $\rm D_{e}=v_{0}^{2}/(2\nu_{r})$ is the effective diffusivity and $\rm v_{0}$ is the effective self-propulsion speed. In the absence of flow and interactions, $\rm v_{0}=v_{p}$ [34]. When interactions are introduced, both $\rm v_{0}$ and $D_{e}$ decrease with increasing packing fraction $\phi$ [14].
Figure 2 shows the MSD of the ABPs with the background flow for various packing fractions $\phi\in[0.1,0.9]$. For low to intermediate packing fractions $\rm\phi\leq 0.5$, the system exhibits the characteristic ballistic-to-diffusive crossover typical of dilute active suspensions [16, 7]. As the system approaches the FIPS regime in high packing fraction limit ($\phi\geq 0.6$), the MSD retains the overall ballistic-to-diffusive transition but develops an additional intermediate “bump” region, indicative of transient trapping in one of the four vortices [20, 35]. A similar MSD bump in non-interacting ABPs [35] originates from particle trapping within the roll-mill vortices.
In Fig. 2, we also observe that the transition time $T_{\rm trans}$ from the ballistic to the diffusive regime is on the order of the fluid timescale, i.e., $2\pi/\omega\sim{\cal O}(10)$. This value is consistent with that reported previously for the FIPS regime [20]. The agreement suggests that, regardless of packing fraction or the emergence of FIPS, the flow predominantly governs the dynamics of the ABPs—a conclusion further supported by the behavior of $\rm v_{0}$ and $\rm D_{e}$ discussed below.
[…]
Figure 2: Mean square displacement (MSD) of ABPs as a function of time for various packing fractions $\phi\in[0.1,0.9]$. The black solid (dashed) lines represent slopes of 2 (1), corresponding to ballistic (diffusive) regimes, respectively, and serve as visual guides in the log-log plot. At higher packing fractions, the emergence of FIPS alters the crossover behavior, leading to deviations from simple ballistic-diffusive dynamics due to persistent clustering. (Top-left inset) normalized Drift velocity $\rm\overline{v}_{d}$ and (bottom-right inset) normalized diffusivity $\rm D_{e}/D_{0}$ as functions of the packing fraction $\rm\phi$. The green curve shows a fit of the form $\rm D_{e}=D_{0}(1-\lambda\phi)^{2}$, with fitting parameter $\rm\lambda=0.8$, indicating a quadratic decay of diffusivity with packing fraction.
(a)Refer to caption(b)Refer to caption(c)
Figure 3: Number fluctuations in ABPs with and without flow, color code is same as Fig.2. The standard deviation $\rm\Delta N_{\ell}$ of particle number is plotted against the mean number $\rm N_{\ell}$ in subregions of size $\rm\ell^{2}$ of ABPs with size $\rm a=0.02$ and scaled speed $\rm V=0.4$, for packing fractions $\rm\phi\in[0.1,0.9]$. (a) In the absence of flow, a transition from normal ($\rm\Delta N_{\ell}\sim N_{\ell}^{1/2}$) to giant fluctuations ($\rm\Delta N_{\ell}\sim N_{\ell}$) is observed as $\rm\phi$ increases beyond 0.4, consistent with the onset of MIPS. (b) In the presence of background flow, a similar crossover is observed near $\rm\phi\approx 0.5$, indicative of FIPS. Dashed and dash-dotted lines serve as visual guides for $\rm\alpha=0.5$ and $\rm\alpha=1$, respectively. (c) Scaling exponent $\rm\alpha$ of the number fluctuations $\rm\Delta N_{\ell}\sim N_{\ell}^{\alpha}$ plotted against packing fraction $\rm\phi$ without (red triangles) and with (blue squares) background flow.
III.2.1 Drift Velocity ($\rm v_{d}$) and Diffusivity ($\rm D_{e}$)
@@
From the MSD curves, we extract both the drift velocity, $\mathrm{v_{d}}$, and the effective diffusivity, $\mathrm{D_{e}}$. The drift velocity for ABPs has been normalized by the self-propulsion speed $\rm v_{p}$} and the root-mean-square ($\rm u_{rms}$) of the background flow $\rm\bm{u(r},t\bm{)}$ for the case without and with flow, respectively. Our results indicate that the normalized drift velocity ($\rm\overline{v}_{d}$) is primarily driven by the background flow, which corresponds to $\mathrm{U_{rms}}=\mathrm{U_{0}}/\sqrt{2}$, as shown in the upper inset of Fig. 2 in blue, for all values of $\phi$. This demonstrates that the externally imposed flow speed dominates over the particle propulsion speed, $\mathrm{v_{p}}$. In contrast, in the absence of background flow, the normalized drift velocity $\mathrm{\overline{v}_{d}}$ becomes a function of packing fraction $\phi$, as reported by Fily et al. [14]. Specifically, the normalized effective self-propulsion speed follows $\mathrm{v_{e}}(\phi)=v_{0}(1-\lambda\phi)$ with $\lambda=0.9$ (shown in lower inset with dotted green line), also shown in the upper inset of Fig. 2 in red color for comparison. In this case, we have normalized $\mathrm{\overline{v}_{d}}$ by the propulsion speed $\mathrm{v_{p}}$. This suggests that self-propulsion and flow together maintain a sustained translation irrespective of crowding, whereas for MIPS there is a strong dependence on crowding.
Interestingly, we observe that the normalized effective diffusivity $\rm\frac{D_{e}}{D_{0}}$ decreases with increasing packing fraction $\rm\phi$, consistent with previous observations in MIPS systems [14, 19]. As shown in the lower inset of Fig. 2, the ratio of effective to absolute diffusivity, $\rm D_{e}/D_{0}$, steadily decreases as the packing fraction $\rm\phi$ increases. This indicates that particles move less freely at higher densities because they are more crowded and interact more often, which slows their overall motion.
The absolute diffusivity $\rm D_{0}$ in our system is defined as:
$\rm D_{0}=\frac{(v_{p}+U_{0})^{2}}{2(\nu_{r}+\omega_{0})},$ (6)
Substituting these values yields $\rm D_{0}\approx 0.487\sim 0.5$.
The observed scaling of the effective diffusivity with packing fraction is well described by the fit[14]:
$\rm D_{e}=D_{0}(1-\lambda\phi)^{2},$ (7)
with a fitting parameter $\lambda=0.8$, indicating a quadratic decay of diffusivity with density as shown in the lower inset with a solid green line.
(a)Refer to caption(b)
Figure 4: Normalized cluster size distribution $\rm P(n)/P(1)$ as a function of ratio of cluster size to maximum number of particle at corresponding packing fractions $\rm n/N_{max}$ for ABPs with size $\rm a=0.02$ at various packing fractions $\rm\phi\in[0.1,0.9]$: (a) without background flow exhibiting MIPS behavior, and (b) with background flow exhibiting FIPS behavior.
(a)Refer to caption(b)
Figure 5: Plot of (a)the exponent $\beta$ and (b) maximum cluster size ($\rm n_{max}$) as a function of packing fraction $\phi$ for ABPs with particle size $a=0.02$. Red triangles (blue diamonds) represent simulations without (with) background flow.
III.3 Number Fluctuations

### 14 Mar 2026 v2

@@ Abstract
We investigate the collective dynamics of active Brownian particles (ABPs) subjected to a steady two-dimensional four-roll-mill flow using numerical simulations. By varying the packing fraction ($\phi$), we uncover a novel flow-induced phase separation (FIPS) that emerges beyond a critical density ($\rm\phi_{c}\gtrsim 0.5$). The mean-square displacement (MSD) exhibits an intermediate plateau between ballistic and diffusive regimes, indicating transient trapping and flow-guided clustering. The effective diffusivity decreases quadratically with $\phi$, while the drift velocity remains nearly constant, demonstrating that large-scale transport is primarily dictated by the background flow. Number fluctuations show a crossover from normal to giant scaling, signaling the onset of long-range density inhomogeneities in the FIPS regime. Our findings provide new insights into the coupling between activity, crowding, and flow, offering a unified framework for understanding phase behavior in driven active matter systems.
@@ I Introduction
The study of self-propelled particle systems gained significant attention following the pioneering work of Vicsek et al. [35], which introduced a minimal model exhibiting collective motion through local alignment interactions. This line of research revealed that the motion of active particles is not only governed by their internal propulsion mechanisms, but is also strongly influenced by the surrounding environment, including background flow fields, confinement, or hydrodynamic interactions [33, 32, 21, 14]. With the rapid advancement of computational resources in recent years, active matter has become a thriving area of research offering deep insights into nonequilibrium statistical mechanics.
@@
While the active Brownian particle (ABP) model neglects hydrodynamic, it has proven to be an indispensable minimal model that captures many features of active matter [6, 1, 31]. In ABPs, each particle self-propels at a constant speed and undergoes rotational diffusion, breaking detailed balance, and giving rise to non-equilibrium behavior even in simple settings [6, 20]. In real-world scenarios, active particles are not isolated but are often immersed in complex fluid environments or interact with background flows, due to external fields, confinement [39, 12, 26]. The interaction between active particles and background flows can give rise to complex dynamical behaviors that are not captured only by considering the ABP model in the absence of external flow influences.
In our recent work [20], we studied ABPs in a two-dimensional steady flow field and discovered a new non-equilibrium phase, which we term flow-induced phase separation (FIPS), driven by the ratio of persistence length of ABPs and background flow. In this study, we investigate the onset and characteristics of the FIPS phase by systematically varying the packing fraction of ABPs in a two-dimensional background flow [20]. As MIPS is known to be strongly dependent on packing fraction [6, 24, 28], we explore FIPS by varying the packing fractions, which range from $0.1$ to $0.9$ to understand how packing fraction influences clustering and phase behavior under flow influence. To characterize the system’s behavior across this range of packing fraction, we compute several key observables: number fluctuations[6, 7, 10, 15], mean squared displacement (MSD) [20, 6, 8], overlap function $\rm Q(t)$ [13, 11, 9], cluster size distributions (CSD) [5, 20], and Probability distribution function (PDF) of Okubo-Weiss parameter ($\Lambda$) [19, 18].
@@ II Model and Simulations
This flow field yields a periodic vorticity pattern in the square simulation domain of size $\rm[0\penalty 10000\ L]^{2}$, with $\rm L=2\pi$, consisting of a pair of clockwise and counterclockwise vortices, separated by narrow regions dominated by strain [20, 34]. This flow has been investigated both experimentally and theoretically in the context of passive tracer mixing [29, 30]. We have set the amplitude of the velocity field to $\rm U_{0}=0.5$.
We use the Euler–Maruyama method to numerically integrate equations (1a) and (1b). The parameters are fixed as $\rm\mu_{t}=1$, $\rm k=100$, and $\rm v_{p}=0.2$. Based on the characteristic velocity scales of the ABPs and the background flow, we define a dimensionless scaled speed $\rm V=\frac{v_{p}}{U_{rms}}$}, where $\rm U_{rms}=U_{0}/\sqrt{2}$ [20]. A characteristic timescale for ABPs is given by $\rm\nu_{r}^{-1}$, while that for the background flow can be defined from the characteristic vorticity $\rm\omega_{rms}=U_{0}$. The dimensionless scaled time is then defined as the ratio of these two timescales, $\rm\tau=\omega_{rms}/(2\pi\nu_{r})$. In all simulations, we set $\rm V\sim 0.5$} and $\tau\sim 25$. Based on our experience from prior work, we have chosen this particular $V$ and $\tau$ such that for $\phi=0.7$, we will get a flow-induced phase separation (FIPS), where ABPs try to phase separate in the strain-dominated region of the flow [20]. We neglect the back-reaction of particles on the background flow, i.e., the flow field remains unaffected by the presence of ABPs [20]. The particle radius is fixed at $\rm a=0.02$ (except in Fig. 3(a) and (b) insets). The packing fraction is defined as $\rm\phi=\frac{N\pi a^{2}}{L^{2}}$ and is varied from $0.1$ to $0.8$ by changing the number of particles $\rm N$ (e.g., $\rm N=21991$ for $\phi=0.7$). All simulations were performed for $10^{4}$ time units, which is sufficient to fully capture the diffusive regime. In all simulations, $\phi$ serves as the control parameter with $\phi\in[0.1,0.9]$.
@@ III Results
In the following subsections, we first present a qualitative comparison between MIPS and FIPS across different values of $\rm\phi$. We then analyze the dynamical behavior of the system using the mean-square displacement (MSD) to estimate the drift velocity and effective diffusivity. The relaxation dynamics are further examined through the overlap function. To quantify structural heterogeneity, we analyze number fluctuations and cluster-size distributions. Finally, to characterize the influence of flow topology on FIPS, we examine the distribution of the Okubo–Weiss parameter.
@@ III.1 Flow-Induced Phase Separation
At high packing fraction ($\phi=0.7$), both systems exhibit phase separation, though their morphologies differ qualitatively. Without flow, we observe standard MIPS, where dense clusters form spontaneously and diffuse throughout the domain (Fig. 1c). In contrast, under imposed flow, the dense phase becomes localized in the strain-dominated regions, forming a characteristic four-lobe structure corresponding to the vortex cores of the four-roll mill flow (Fig. 1f). A systematic variation of $\phi$ from $\rm\phi\in[0.1,0.9]$ is shown in Appendix Fig. 6, demonstrating that even at very high packing fractions ($\phi=0.9$), the four distinct low-density regions persist, coinciding with the four vortex cores of the background flow.
In this system, $\rm\mathcal{O}(v_{p})=\mathcal{O}(U_{0})$, while the flow field is spatially varying. Consequently, in regions where $\rm|\mathbf{u}|>v_{p}$, particle motion is primarily governed by the external flow, which typically corresponds to regions of low vorticity, causing particles to follow the streamlines. In contrast, in regions where $\rm|\mathbf{u}|<v_{p}$, the dynamics are dominated by the self-propulsion velocity of the particles. These regions are associated with larger vorticity, which strongly influences the particle orientation and direction of motion. At low packing fractions, particles have sufficient free space to explore the domain, resulting in a nearly homogeneous spatial distribution (Fig. 1(d) and 1(e)). However, at high packing fractions, interparticle interactions significantly affect the dynamics. Starting from a homogeneous configuration, particles in the strain-dominated regions largely follow the background flow, whereas particles within vortices experience strong vorticity, enabling them to escape the vortices and accumulate in the strain-dominated regions. This process ultimately leads to FIPS (Fig. 1(f)).
(a)(b)(c)(d)
Figure 2: Dynamical behavior of ABPs under background flow. (a) MSD normalized by the domain size as a function of time normalized by the flow timescale ($\rm t_{f}$) for packing fractions $\rm\phi=0.1$$0.8$. A ballistic–diffusive crossover is observed, with an intermediate plateau at large $\rm\phi$ due to particle trapping in strain-dominated regions and near vortex peripheries, away from the cores, as illustrated by the trajectories in (d). Insets: Normalized drift velocity (top-left) and effective diffusivity (bottom-right) for cases with (squares) and without (asterisks) flow, obtained from MSD fits as functions of $\rm\phi$. (b) Overlap function $\rm Q(t)$ versus time showing delayed relaxation at high packing fractions and an intermediate modulation associated with transient trapping in strain-dominated regions and vortex peripheries. The black dotted line in (a) and (b) corresponds to non-interacting ABPs in background flow. Inset: Normalized relaxation time versus $\rm\phi$ for FIPS and MIPS defined as $\rm Q(t)=1/e$. While both increase exponentially with $\rm\phi$, MIPS shows a pronounced rise near $\rm\phi_{c}=0.3$. The relaxation time is normalized by the value in the non-interacting case. (c) and (d) Particle trajectories at $\rm\phi=0.1$ and $0.8$, respectively, at simulation time $\rm t=1000$.
@@ III.2 Mean Square Displacement (MSD)
$\rm\langle[\Delta\boldsymbol{r}(t)]^{2}\rangle=4D_{0}\left[t+\frac{1}{\nu_{r}}(e^{-\nu_{r}t}-1)\right],$ (5)
@@
where $\rm D_{0}=v_{p}^{2}/(2\nu_{r})$ is the absolute diffusivity and $\rm v_{p}$ is the absolute self-propulsion speed [25]. When interactions are introduced, overall MSD decreases with increasing packing fraction $\phi$ [6].
Fig. 2(a) shows the mean-square displacement (MSD), normalized by the domain size, for ABPs in the presence of background flow at different packing fractions $\phi\in[0.1,0.9]$. For low to intermediate packing fractions $\rm\phi\leq 0.5$, the system exhibits a characteristic ballistic-to-diffusive crossover approximately at fluid time scale $\rm t_{f}=\rm 2\pi/\rm\omega_{rms}$. This can be understood by examining the trajectories shown in Fig. 2(c). At short times $\rm t\ll t_{f}$, the particles move along nearly persistent paths, leading to ballistic scaling $\rm\langle[\Delta\boldsymbol{r}(t)]^{2}\rangle\sim t^{2}$. As $\rm t$ approaches the fluid time scale $\rm t_{f}$, the trajectories curve and lose directional correlation. For $\rm t\gtrsim t_{f}$, the motion becomes effectively randomized, resulting in diffusive behavior $\rm\langle[\Delta\boldsymbol{r}(t)]^{2}\rangle\sim t$. As the system approaches the FIPS regime in the high packing fraction limit ($\rm\phi\geq 0.6$), the MSD retains the overall ballistic-to-diffusive transition but develops an additional intermediate “plateau” region [20]. In Fig. 2(d), at short times $\rm t\ll t_{f}$, particles exhibit nearly persistent motion, resulting in ballistic scaling. As $\rm t$ approaches the fluid time scale $\rm t_{f}$, particles become temporarily trapped between strain-dominated regions and near vortex peripheries, away from the vertex cores due to crowding, leading to a plateau in the MSD. Eventually, they escape these cages and exhibit diffusive motion for $\rm t\gtrsim t_{f}$. In Fig. 2(a), the value $\rm\mathrm{MSD/L^{2}}\big|_{\mathrm{t=t_{f}}}\sim\rm 0.1-0.2$ is consistent with the expected vortex size, which is approximately $\rm L/4-L/2$. To highlight the role of interactions, we additionally plotted the MSD of non-interacting ABPs in a background flow, indicated by the black dotted line in Fig. 2(a), which shows good agreement at low packing fractions.
Drift Velocity $\rm(v_{d})$ and Diffusivity $\rm(D_{e})$: From the MSD curves, we determine both the drift velocity, $\mathrm{v_{d}}$, and the effective diffusivity, $\mathrm{D_{e}}$. The drift velocity of the ABPs is normalized by a characteristic velocity scale obtained from the non-interacting case. In the absence of flow, this scale corresponds to the self-propulsion speed $\rm v_{p}$}, whereas in the presence of flow it is given by the characteristic velocity, i.e., root-mean-square velocity, $\mathrm{U_{rms}}$, of the background flow field $\rm\mathbf{u}(\mathbf{r},t)$. In the absence of background flow, the normalized drift velocity $\rm\overline{v}_{d}$ decreases with packing fraction $\phi$. The effective propulsion speed follows $\rm v_{d}(\phi)=v_{p}(1-\lambda\phi)$ with $\lambda=0.9$, shown in the upper inset of Fig. 2 (red asterisk) [6].
@@
In the presence of the flow, the $\rm\overline{v}_{d}$ is nearly constant which indicate that the normalized drift velocity ($\rm\overline{v}_{d}$) is primarily driven by the background flow, which corresponds to $\mathrm{U_{rms}}=\mathrm{U_{0}}/\sqrt{2}$, as shown in the upper inset of Fig. 2(a) in blue square, for all values of $\phi$. This suggests that self-propulsion and flow together maintain a sustained translation irrespective of crowding, whereas for MIPS, there is a strong dependence on crowding.
Similarly, the extracted effective diffusivity is normalized by the absolute diffusivity of the corresponding non-interacting system. For the MIPS case, this diffusivity is given by $\rm D_{0}=v_{p}^{2}/2\nu_{r}$ [6]. In contrast, for the FIPS case we obtain $\rm D_{0}\sim 0.5$. Determining the exact analytical form of the absolute diffusivity in the presence of flow is beyond the scope of the present work and will be addressed in future studies.
As shown in the lower inset of Fig. 2(a), the normalized effective diffusivity $\rm D_{e}/D_{0}$ decreases with increasing packing fraction $\rm\phi$. In the MIPS case, this decrease exhibits a sharp transition near the critical packing fraction $\rm\phi_{c}\sim 0.3$, consistent with earlier reports [16]. In contrast, for FIPS, $\rm D_{e}/D_{0}$ decreases smoothly with $\rm\phi$, indicating a continuous transition. For MIPS, the data are well described by $\rm D_{e}=D_{0}(1-\lambda_{1}\phi^{2})$ for $\rm\phi<0.3$ (yellow plus markers), with $\rm\lambda_{1}=1$. For $\rm\phi>0.3$, the behavior is better captured by $\rm D_{e}=D_{0}\lambda_{2}(1-\lambda_{1}\phi)^{2}$, with $\rm\lambda_{2}=0.4$ (green triangles), as reported by Nayak et al. [16]. In the FIPS case, the diffusivity follows a different scaling, $\rm D_{e}=D_{0}(1-\lambda_{3}\phi)$, with $\rm\lambda_{3}=0.8$. This indicates that particles move less freely at higher densities because they are more crowded and interact more frequently, slowing their overall motion.
III.3 Overlap Function
Fig 2(b) is the typical curve of time-dependent self overlap function $\rm Q(t)$ for different $\rm\phi$ is defined as
$\mathrm{Q(t)=\frac{1}{N}\langle\sum_{j=1}^{N}}w(|\mathrm{\mathbf{r}_{j}(t)-\mathbf{r}_{j}(0)}|)\rangle,$ (6)
where $\rm N$ denotes the number of particles and the weight function $w(\mathrm{x})$ is defined as $w(\mathrm{x})=1$ for $\rm x<\pi$ (i.e. within the vortex size) and $w(\mathrm{x})=0$ otherwise [13],[9].
Fig 2(b) shows $\rm Q(t)$ exhibits an approximately exponential decay with some modulations depends on $\rm\phi$. At low packing fractions, particles escape their cages before the flow time scale $\rm t_{f}$, indicating that the decorrelation dynamics are governed by both the flow and the intrinsic activity of the particles. The self-propulsion enhances particle motion and consequently reduces the cage escape time. Around $\rm t=t_{f}$, a small modulation in the $\rm Q(t)$ curve suggests that the flow tends to recage the particles due to its periodic nature; however, the persistent activity prevents sustained recaging.
After the critical packing fraction $\rm\phi_{c}\gtrsim 0.5$ for FIPS, particle interactions become dominant, which delays the cage-out process. Near $\rm t=t_{f}$, recaging is enhanced due to crowding between vortices and strain-dominated regions. The clusters as a whole become trapped in and around the strain-dominated regions, as indicated by the pronounced modulation in the $\rm Q(t)$ curve. Eventually, particles escape the cage due to their self-propelled velocity.
The $\rm Q(t)$ plot shows stronger caging and recaging at high packing fractions, indicating that particles remain trapped for longer times before escaping, consistent with the plateau observed in the MSD. To highlight the role of interactions, we also plot $\rm Q(t)$ for non-interacting ABPs in the same flow field, which shows good agreement with the interacting case at low packing fractions.
Relaxation time: The relaxation time, $\rm\tau_{e}$, is defined as the time at which the correlation function satisfies $\rm Q(t)=1/e$, where $\rm e$ is the base of the natural logarithm [9]. The inset of Fig. 2(b) shows the normalized relaxation time, $\rm\tau_{e}/\tau_{0}$, as a function of the packing fraction $\rm\phi$ for both without and with background flow cases, where $\rm\tau_{0}$ denotes the relaxation time in the non-interacting case.
For both cases, the relaxation time is growing exponentially as $\rm\phi$ increases which means basic slowdown mechanism is interaction-driven in both cases. The system without flow exhibits a relatively sharp increase around $\rm\phi\approx 0.3$, indicating a sudden onset of clustering due to activity-driven phase separation. In contrast, the FIPS case shows a smoother variation of the relaxation time, suggesting that the background flow continuously redistributes particles and leads to a more gradual buildup of caging and clustering.
(a)(b)(c)
Figure 3: Number fluctuations in ABPs varying with packing fraction for (a) system without background flow shows giant number fluctuations for $\phi\geq 0.36$ and (b) system with flow shows giant number fluctuations for $\phi\gtrsim 0.5$. Insets in (a) and (b) show the scaling exponent $\alpha$ of the number fluctuations as a function of packing fraction $\phi$ for different system sizes, $\rm N=880$$10^{4}$, providing a good statistics for large number of particles for both without and with flow cases. (c) A direct comparison between without and with flow cases exhibits a sharp and smooth transition, respectively.
III.4 Number Fluctuations